Shift report 20000813 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sun Aug 13 2000 - 00:00:55 EDT

  • Next message: http account: "Shift report 20000813 00:00-08:00"

    ------------------------------------------------------------
    Date: Sun Aug 13 00:00:54 2000, Shift: 16:00-24:00
    Supervisor: HB, with: HB & CE
    Log book#4 pages 74 - 76
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam on from approx 21.40. Prepare to dump beam 
       from 23.30. 
    2. Two runs. run002232 (no voltage on ZDCs) + 
       run002233 (good - MRS at 45, D5 off).
       They called from the Main Control Room and said
       that they didn't get any ZDC data from us. It looked
       fine on our "RHIC" monitor (after we turned on the HV).
    3. Nothing to report.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sun Aug 13 2000 - 00:04:30 EDT